# Interpretable Multi-task Learning with Shared Variable Embeddings

### v1

@@ 4.1 Classification capability
4.2 Interpretability
  • Compute the attention scores for all raw variable embeddings.
The variables most similar to a shared embedding representing a specific concept would be expected to share some intuitive notion, category or semantic meaning. In our evaluations, we choose $K=5$. To aid in a quantitative as well as qualitative assessment of mapping to concepts, we introduce a measurable and less subjective assignment to real-world concepts in the form of the Subject Area ascribed to a given task/dataset in the UCI repository [20]. Each dataset has one Subject Area (SA) assigned to it from the following $11$ possibilities: Biology (Bio), Business (Bus), Climate and Environment (C&E), Computer Science (CS), Engineering (E), Games (G), Health and Medicine (H&M), Law (L), Physics and Chemistry (P&C), Social Sciences (SS) and Other (O). These categories afford us the option to measure the performance of a given model in terms of interpretability and compare it with other methods.
@@ 4.2 Interpretability
Table 4 and Table 5 show the results of running our evaluation procedure for one randomly chosen shared embedding. This consists of $5$ variables most similar to the sampled shared embedding. We also present extended results where this procedure is repeated in Appendix A.
Table 5 shows a similar result for the $1.05$-entmax sparse attention method. In this case, the majority of the selected variables seem to represent concepts related to health or biological systems. In quantitative terms, the majority belongs to one Subject Area - Health and Medicine, while there is also one variable identified as coming from the Biology Subject Area. The one variable which does not fit the health or biological interpretation is the least similar from the selected $5$ and it belongs to the by-definition-broad Other category. Qualitatively, the most similar variables relate to living organisms. This is also the case for the fourth most similar variable from the Biology SA. We do, however, notice less internal consistency in this grouping, relative to the results from the base method. Both quantitative and qualitative results suggest that the base shared embedding method may actually produce more interpretable shared embeddings than the sparse attention approach. This is supported by repeat analysis presented in Appendix A where both methods seem to produce interpretable embeddings but the base method outperforms the sparse attention method in terms of the cohesion of the embeddings. The base method produces representations which are more easily linked to one broad intuitive concept while the inclusion of sparse attention prefers embeddings which are linked to more than one but still related concepts (e.g. biological and health-related ones).
While this investigation points to the relative performance of the methods, it is worth analyzing whether the results are not merely caused by the statistical characteristics of the dataset. To facilitate this, we replace the evaluation procedure with its random counterpart where we sample $K=5$ variables from the UCI-121 dataset and perform the same qualitative and quantitative assessment as was the case for variables similar to the shared embedding from our models. The results for one such sample are presented in Table 6. There is one repeated category (H&M), however, it does not have a majority. Also, other than in some samples for the $1.05$-entmax method, there is only one repeat category, not a contest between two categories. These indications also hold across additional samples presented in Appendix A. A further differentiating characteristic is that all the variables within each of the extended samples (in Appendix A) come from different datasets, whereas for our analyzed models datasets occasionally repeat. The only case when we observe a majority category for the random variables is one where these variables represent the Other SA, which is, by definition, a broad bracket in which we do not expect the variables to represent similar concepts. All this points to the fact that the most similar variables obtained from out models are significantly different from random choice.
4.3 Accuracy vs. interpretability trade-off
Drawing on the quantitative and qualitative results, we find that for our best performing methods none of them strictly dominates the other in terms of both accuracy and interpretability. The $1.05$-entmax method achieves higher final accuracy than our base shared embedding method, $81.9\%$ vs. $81.5\%$. On the flip side, the base shared embedding method does seem to more successfully separate real-world concepts into specific shared embeddings. Specifically, in our extended results (Appendix A), we see that the base shared embedding method achieves more consistent concept assignment to the shared embeddings. For $5$ trials, we obtain $3$ where there is a majority SA. In one trial, there was no majority but there was a dominant SA without draws. One trial resulted in a draw between SAs. It is also worth noting that for one trial, all the similar variables come from the same SA (P&C) and from different datasets. Conversely, for the $1.05$-entmax method, the assignment of shared embeddings to concepts is still present, only weaker. One trial results in a majority SA assignment (H&M). $3$ trials end in draws between dominant categories. Importantly, one trial has all the similar variables represent different SAs from distinct datasets. With these results, there seems to be a trade-off between prediction accuracy and interpretability.
@@ 5 Conclusion
In empirical experiments, we have shown that our base method performs as well as the standard variable embedding method on the UCI-121 dataset, while not requiring any fine-tuning, which the standard method does. Additionally, we have performed a series of ablations to identify which versions of our architecture perform favorably in terms of classification accuracy and the potential interpretability of the shared embeddings. The results have demonstrated that the sparse attention mechanism helps in: (1) achieving superior classification performance and (2) requiring significantly less training steps than our base method. However, the gain comes at a cost of decreased interpretability relative to our base shared embedding method. This suggests a potential trade-off between performance and interpretability.
As far as interpretability itself is concerned, both our base method and its extension with sparse attention are able to use the shared embeddings to identify abstract concepts instead of making hard links to concrete variables from specific datasets, which is the case for the standard variable embedding approach. The base shared embedding method generates embeddings which are more interpretable and internaly consistent than the sparse attention modification.
@@
Appendix A Extended interpretability results.
While the samples presented in the main paper are instructive of the ability of our models to produce interpretable shared embeddings and of the difference between them and a random assignment, it is important to present and analyze a larger number of samples. Additional samples for (a) the base shared embedding method, (b) the $1.05$-entmax sparse attention method and (c) random choice are presented below in Tables 7, 8, 9. For each method, the analysis encompasses 5 trials. In each trial, a shared embedding is chosen at random without replacement. For the base shared embedding method and the $1.05$-entmax method, $5$ variables most similar to the chosen random shared embedding are presented. The similarity between the sampled shared embedding $\mathbf{s}_{p}$ and the $i$-th variable is measured as the cosine similarity $S_{C}$ between the shared embedding and the processed variable embedding $\mathbf{f}_{i}$ associated with this specific variable:
@@
A.1 Shared embedding method
@@
    There is a dominant category with $4$ variables in the form of Health and Medicine. All the selected variables come from different datasets. There is significant coherence to the grouping, which can also be seen in qualitative terms as the only physical variable in the selection can still be understood as describing elements of a real-world structure, similar to most of the biological variables. All in all, an intuitive biological concept can be identified.
@@
    There is a dominant category, Health and Medicine, albeit not a majority category. There are $4$ distinct SA represented and $4$ datasets. The majority of the selected variables can still intuitively be interpreted as ones related to health or the biological functioning of organisms, but outliers such as values from synthetically generated charts are also present. Overall, the interpretation is made significantly harder by ambiguous data.
@@
    We do have a dominant SA, Biology, with $2$ variables, but again, there is another category with the same number of identified variables - Health and Medicine. Also, we see that for this trial, there are 2 repeated datasets. The intuitive meaning behind the variables from this trial can be interpreted as describing living things but more details are ocluded by the fact that all the variables for the Health and Medicine SA are ambiguous.
@@ A.3 Random choice
In order to account for the statistical properties of the UCI-121 dataset, we perform an analysis where the selected variables are actually randomly sampled without repetition from the dataset. If the dataset is not heavily skewed toward the concepts identified by either of our methods, it is natural to assume that we will see a lot more variability in the selection. For a random choice of variables, one could expect not to see majority SAs, or at least see them infrequently. Similarly, the expectation would be to see more SAs within each trial than is the case for our methods. Also, a random assignment would result in very frequent situations where all the variables come from distinct datasets. Conversely, for the base shared embedding method and the $1.05$-entmax method the expectation would be that the variables most similar to a given shared embedding would be more likely to come from the same dataset. Table 9 summarizes the results for the random choice of variables. Indeed, there is only one trail with a majority category, but on inspection the identified SA is Other, which is a blanket category for a range of datasets representing different concepts. Apart from this special case, there are no other majority categories in trials. This suggests significantly weaker interpretability than for the base shared embedding method. The sparse attention method does show similar levels of dominant categories, however, with a crucial distinction. In the sparse attention approach, all the non-majority cases bar one had a tie for the dominant category, suggesting that the method was able to identify concepts better than random choice, with the assignment to two competing concepts. For random choice, such a competing assignment does not occur. Also, there are visibly more SAs represented than for the shared embedding method. Qualitatively, random choice does result in an assortment of more than two distinct concepts for a given trail rather than in the identification of an abstract notion or two such notions, which is a frequent situation for the shared embedding method and $1.05$-entmax methods.
Appendix B Architecture of the proposed method.

### v2

@@ 4.2 Interpretability
Table 4 and Table 5 show the results of running our evaluation procedure for one randomly chosen shared embedding. This consists of $5$ variables most similar to the sampled shared embedding. We also present extended results where this procedure is repeated in Appendix B and results on SVEs commonly shared by tasks in Appendix G.
@@
Table 5 shows a similar result for the $1.05$-entmax sparse attention method. In this case, the majority of the selected variables seem to represent concepts related to health or biological systems. In quantitative terms, the majority belongs to one Subject Area - Health and Medicine, while there is also one variable identified as coming from the Biology Subject Area. The one variable which does not fit the health or biological interpretation is the least similar from the selected $5$ and it belongs to the by-definition-broad Other category. Qualitatively, the most similar variables relate to living organisms. This is also the case for the fourth most similar variable from the Biology SA. We do, however, notice less internal consistency in this grouping, relative to the results from the base SVE method. Both quantitative and qualitative results suggest that the base shared embedding method may actually produce more interpretable shared embeddings than the sparse attention approach. This is supported by repeat analysis presented in Appendix B where both methods seem to produce interpretable embeddings but the base method outperforms the sparse attention method in terms of the cohesion of the embeddings. The base method produces representations which are more easily linked to one broad intuitive concept while the inclusion of sparse attention prefers embeddings which are linked to more than one but still related concepts (e.g. biological and health-related ones).
@@
While this investigation points to the relative performance of the methods, it is worth analyzing whether the results are not merely caused by the statistical characteristics of the dataset. To facilitate this, we replace the evaluation procedure with its random counterpart where we sample $K=5$ variables from the UCI-121 dataset and perform the same qualitative and quantitative assessment as was the case for variables similar to the shared embedding from our models. The results for one such sample are presented in Table 6. There is one repeated category (H&M), however, it does not have a majority. Also, other than in some samples for the $1.05$-entmax method, there is only one repeated category, not a contest between two categories. These indications also hold across additional samples presented in Appendix B. A further differentiating characteristic is that all the variables within each of the extended samples (in Appendix B) come from different datasets, whereas for our analyzed models datasets occasionally repeat. The only case when we observe a majority category for the random variables is one where these variables represent the Other SA, which is, by definition, a broad bracket in which we do not expect the variables to represent similar concepts. All this points to the fact that the most similar variables obtained from out models are significantly different from random choice.
4.3 Accuracy vs. interpretability trade-off
Drawing on the quantitative and qualitative results, we find that for our best performing methods none of them strictly dominates the other in terms of both accuracy and interpretability. The $1.05$-entmax method achieves higher final accuracy than our base shared embedding method, $81.9\%$ vs. $81.5\%$. On the flip side, the base shared embedding method does seem to more successfully separate real-world concepts into specific shared embeddings.
@@
Specifically, in our extended results (Appendix B), we see that the base shared embedding method achieves more consistent concept assignment to the shared embeddings. For $5$ trials, we obtain $3$ where there is a majority SA. In one trial, there was no majority but there was a dominant SA without draws. One trial resulted in a draw between SAs. It is also worth noting that for one trial, all the similar variables come from the same SA (P&C) and from different datasets. Conversely, for the $1.05$-entmax method, the assignment of shared embeddings to concepts is still present, only weaker. One trial results in a majority SA assignment (H&M). $3$ trials end in draws between dominant categories. Importantly, one trial has all the similar variables represent different SAs from distinct datasets. With these results, there seems to be a trade-off between prediction accuracy and interpretability. It should be noted that, even in the presence of such a trade-off, both the base shared embedding method and the sparse attention method more decisively than random choice link specific shared embeddings to intuitive concepts. A discussion on this is presented in Appendix B.3.
@@ 5 Conclusion
In empirical experiments, we have shown that our base method performs as well as the standard variable embedding method on the UCI-121 dataset, while not requiring any fine-tuning, which the standard method does. Additionally, we have performed a series of ablations to identify which versions of our architecture perform favorably in terms of classification accuracy and the potential interpretability of the shared embeddings. The results have demonstrated that the sparse attention mechanism helps in: (1) achieving superior classification performance and (2) requiring significantly less training steps than our base SVE method. However, the gain comes at a cost of decreased interpretability relative to our base shared embedding method. This suggests a potential trade-off between performance and interpretability.
As far as interpretability itself is concerned, both our base method and its extension with sparse attention are able to use the shared embeddings to identify abstract concepts instead of making hard links to concrete variables from specific datasets, which is the case for the standard variable embedding approach. The base shared embedding method generates embeddings which are more interpretable and internally consistent than the sparse attention modification.
@@
Acknowledgements
This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology.
Appendix A Details of methods imposing independence of shared variable embeddings
A.1 Orthogonalization
For a simple notion of independence, we consider an embedding to be independent from other embeddings when it is not a linear combination of them. With this, $r_{\mathbf{S}}=\text{rank}(\mathbf{S})$ is a measure of independence. In a realistic setting, we might still have $r_{\mathbf{S}}=C$ even if multiple embeddings are approximately linearly dependent. An operational measure of the rank of $\mathbf{S}$ would require to address this drawback and we describe such a measure $q_{\mathbf{S}}$ in Section A.2. The results from Section 4.2 show that the proposed training procedure results in $q_{\mathbf{S}}<C$. A straightforward way to build in more independence is to require $\mathbf{S}$ to consist of orthonormal vectors. With an additional assumption of $D=C$, this would translate into an orthogonality requirement, which could be incorporated in the loss function:
A.2 Stable rank
Instead of focusing on restricting $\mathbf{S}$, it is possible to explicitly add $r_{\mathbf{S}}$ to the loss function. A significant drawback of this is the discontinuous characteristic of the rank measure, which makes it unsuitable for gradient descent. To address this, we rely on a continuous proxy. Let us consider a matrix $\mathbf{A}_{N\times M}$ with $\sigma_{i}(\mathbf{A})$ being its $i$-th singular value. The Frobenius norm of $\mathbf{A}$ is defined as $\left\lVert\mathbf{A}\right\rVert_{F}^{2}=\text{tr}(\mathbf{A}\mathbf{A}^{T})=\sum_{i,j}{\mathbf{A}_{i,j}^{2}}=\sum_{i}{\sigma_{i}^{2}}$. The stable rank (Vasershtein 1971) of $\mathbf{A}$ is then defined as:
and $\text{sr}(\mathbf{A})\leq\text{rank}(\mathbf{A})$.
A.3 Von Neumann entropy
$V(\mathbf{A})=-\sum_{i}{\sigma_{i}^{2}\ln{\sigma_{i}^{2}}}$ (11)
A.4 Sparse attention
In a procedure orthogonal to inducing structure in the shared embedding matrix, one can also restrict the way in which the actual shared embeddings are combined to form the processed embeddings. One drawback of the standard attention mechanism is that it assigns non-zero weights to all the value vectors. This means that even components with marginal similarity to the keys are present in the final linear combinations. A potential solution would be to make the output of the attention mechanism not rely on values with small similarity scores. In order to keep the whole mechanism differentiable, we adopt the $\alpha$-entmax method (Peters, Niculae, and Martins 2019). Let us denote the $d$-probability simplex by $\triangle^{d}=\{\mathbf{p}\in\mathbb{R}^{d}:\mathbf{p}\geq 0,\left\lVert p\right\rVert_{1}=1\}$. Sparsemax (Martins and Astudillo 2016) is defined as:
A family of Tsallis $\alpha$-entropies (Tsallis 1988) can be defined for $\alpha\geq 1$ as:
where $H^{S}(\mathbf{p})=-\sum_{j}{p_{j}\ln{p_{j}}}$.
Appendix B Extended interpretability results
While the samples presented in the main paper are instructive of the ability of our models to produce interpretable shared embeddings and of the difference between them and a random assignment, it is important to present and analyze a larger number of samples. Additional samples for (a) the base shared embedding method, (b) the $1.05$-entmax sparse attention method and (c) random choice are presented below in Tables 7, 8, 9, respectively. For each method, the analysis encompasses 5 trials. In each trial, a shared embedding is chosen at random without replacement. For the base shared embedding method and the $1.05$-entmax method, $5$ variables most similar to the chosen random shared embedding are presented. The similarity between the sampled shared embedding $\mathbf{s}_{p}$ and the $i$-th variable is measured as the cosine similarity $S_{C}$ between the shared embedding and the processed variable embedding $\mathbf{f}_{i}$ associated with this specific variable:
@@ Appendix B Extended interpretability results
B.1 Shared embedding method
@@
    There is a dominant category with $4$ variables in the form of Health and Medicine. All the selected variables come from different datasets. There is significant coherence in the grouping, which can also be seen in qualitative terms as the only physical variable in the selection can still be understood as describing elements of a real-world structure, similar to most of the biological variables. All in all, an intuitive biological concept can be identified.
@@
    There is a dominant category, Health and Medicine, albeit not a majority category. There are $4$ distinct SA represented and $4$ datasets. The majority of the selected variables can still intuitively be interpreted as ones related to health or the biological functioning of organisms, but outliers, such as values from synthetically generated charts, are also present. Overall, the interpretation is made significantly harder by ambiguous data.
@@
    We do have a dominant SA, Biology, with $2$ variables, but again, there is another category with the same number of identified variables - Health and Medicine. Also, we see that for this trial, there are 2 repeated datasets. The intuitive meaning behind the variables from this trial can be interpreted as describing living things but more details are occluded by the fact that all the variables for the Health and Medicine SA are ambiguous.
@@ B.3 Random choice
In order to account for the statistical properties of the UCI-121 dataset, we perform an analysis where the selected variables are actually randomly sampled without repetition from the dataset. If the dataset is not heavily skewed toward the concepts identified by either of our methods, it is natural to assume that we will see a lot more variability in the selection. For a random choice of variables, one could expect not to see majority SAs, or at least see them infrequently. Similarly, the expectation would be to see more SAs within each trial than is the case for our methods. Also, a random assignment would result in very frequent situations where all the variables come from distinct datasets. Conversely, for the base shared embedding method and the $1.05$-entmax method the expectation would be that the variables most similar to a given shared embedding would be more likely to come from the same dataset. Table 9 summarizes the results for the random choice of variables. Indeed, there is only one trail with a majority category, but on inspection the identified SA is Other, which is a blanket category for a range of datasets representing different concepts. Apart from this special case, there are no other majority categories in trials. This suggests significantly weaker interpretability than for the base shared embedding method. The sparse attention method does show similar levels of dominant categories, however, with a crucial distinction. In the sparse attention approach, all the non-majority cases bar one had a tie for the dominant category, suggesting that the method was able to identify concepts better than random choice, with the assignment to two competing concepts. Overall, for the sparse attention method, $4/5$ trials either had a majority category or a tied dominant category. For random choice, excluding the kitchen sink Other SA, a majority category or a draw between two competing categories occurs in $2/5$ trials. This suggests that for random choice there is less concentration in SAs. Also, there are visibly more SAs represented than for the shared embedding method. The shared embedding method has an average of $2.4$ SAs per trial. The same metric for random choice stands at $3.2$. Qualitatively, random choice does result in an assortment of more than two distinct concepts for a given trail rather than in the identification of an abstract notion or two such notions, which is a frequent situation for the shared embedding method and $1.05$-entmax methods.
Appendix C Architecture of the proposed method